Eleanor is working her way through school. She works two part-time jobs for a total of 25 hours a week. Job A pays $6.00 per hour and job B pay $6.80 per hour. How many hours did she work at each job the week that she made $159.60?
Answer by stanbon(75887)

Eleanor is working her way through school. She works two part-time jobs for a total of 25 hours a week. Job A pays $6.00 per hour and job B pay $6.80 per hour. How many hours did she work at each job the week that she made $159.60?
----
pay + pay = pay
---
6a + 6.8(25-a) = 159.60
-------------------------------
6a + 170 - 6.8a = 159.60
-----------------
-0.8a = -10.40
----
a = 13 hrs (# of hrs worked at job A)
----
b = 25-a = 12 hrs (# of hrs worked at job B)
-----------
Cheers,
Stan H.
--------------